If Sam took a 45 question test and got 5 wrong, bummer, what percent did he get correct on the test?
Answer by jim_thompson5910(35256)

He got 5 wrong out of 45 total. So he got 45-5 = 40 correct.

(# correct)/(# total) = 40/45 = 8/9 = 0.8889 (this is approximate)

Take 0.889 and multiply it by 100 to convert it to a percentage

0.8889*100 = 88.89%

So he got approximately 88.89% correct on his test.